OpenStudy (anonymous):

If g(x) = 3(x - 2), find the value of x if g(x) = 6. A. 12 B. 2.6 C. 8 D. 4

OpenStudy (anonymous):

if g(x)=3(x-2) and g(x)=6, you can set those equations equal to each other and solve for x. So, 6=3(x-2) and go from there.

OpenStudy (anonymous):

so D ?

OpenStudy (anonymous):

@entrails

OpenStudy (anonymous):

Let f(p) be the average number of days a house stays on the market before being sold for price p in $1,000s. Which statement best describes the meaning of f(250)? The house sold for $250,000. The house stayed on the market for an average of 250 days before being sold. This is the average number of days the house stayed on the market before being sold for $250,000. The house sold on the market for $250,000 and stayed on the market for an average of 250 days before being sold.

OpenStudy (anonymous):

Actually, I'm not seeing the answer in the options...here's my work: \[6=3(x-2)\] \[2=x-2\] \[x=4\]

OpenStudy (anonymous):

Are you sure you input the question correctly?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

d'oh, I'm sorry. Yes, the answer is D, x=4. I didn't seem to see that to begin with.

OpenStudy (anonymous):

And in answer to your second question, the answer would be that the house stayed on the market for 250 days and was sold at a price at $250,000 since p is a function of f

OpenStudy (anonymous):

so D again ??

OpenStudy (anonymous):

@entrails

OpenStudy (anonymous):

No, C

OpenStudy (anonymous):

Because f(p)=average number of days the house stayed on market in terms of p, the price. So f(250)=average number of days the house stayed on the market at $250,000

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

For the function f(x) = x + 4, what is the ordered pair for the point on the graph when x = 3p? A. (x, x + 4) B. (x, 3p + 4) C. (3p, x + 4) D. (3p, 3p + 4)

OpenStudy (anonymous):

@entrails

OpenStudy (anonymous):

Okay, so f(x) is essentially the y-value, so that's the second number in the ordered pair. x is the x-value, so in f(3p), the ordered pair would be (x, f(x)) or (3p, f(3p)). With that knowledge, plug in 3p to your function and see what the ordered pair is.